solve:
A) x=-2y
3x-2y=8
B) 2x-3y=13
3x+y=3
C) 2s-t=4
3s-5t=2
D) x/3-y/6=-1/3
x+y=-3-2x-y

Answer by midwood_trail(310)

I will do (A) only.
Follow my steps and do the rest.
QUESTION (A)
x = -2y.......Equation A
3x - 2y = 8...Equation B
Notice how conveniently Equation A already gives us the fact that x = -2y.
Replace x in Equation B with -2y to find the value of y.
3(-2y) - 2y = 8
-6y - 2y = 8
-8y = 8
y = 8/-8
y = -1
I just found the value of y to be -1.
To find the value of x, replace y in EITHER EQUATION (your choice) with -1 and simplify.
I will use Equation A because it is easier to play with.
x = -2y...Equation A
x = -2(-1)
x = 2
The solution for this system of linear equations is the point (2, -1).
What does that point really mean anyway?
The point (2, -1) is the location on the xy-plane where the two given lines meet or cross each other.
Understand?